# Sloth: scaling laws for LLM skills to predict multi-benchmark performance across families

### v3

@@ 4.2 Comparing scaling laws in terms of prediction errors
Refer to caption Refer to caption
Figure 2: Average prediction error across LLM families and benchmarks. Sloth performs well while the scaling laws in which intercept and slope are family-dependent underperform.
@@
In the first set of experiments, we consider the case in which only the smallest model of the test family is observed at training time. Because of that reason, we cannot fit the general scaling law in (2.2) in which both the intercept and slope are family dependent. In this scenario, we consider our main baselines to be (i) the model in (2.2) with shared intercept and slope (Owen, 2024) (“FLOPs (shared intercept)”), (ii) the same model but only with shared slope (“FLOPs”), (iii) a version of the PCA idea [^2] We include more details about the PCA approach in Appendix E. proposed by Ruan et al. (2024) in which we predict the principal components using the FLOPs model with shared slope that are then mapped to the benchmark scores (“PCA + FLOPs”), (iv) and our model with trainable activation function but assuming $\Lambda$ is identity (“Size and Tokens”; implies $d=J$). Moreover, we include two versions of Sloth. In the “basic” one, we fix $\sigma$ to be sigmoid, and $\gamma_{j}$’s are given by the $100\%$ over the number of alternatives in the case of multiple-choice benchmarks [^3] When the benchmark has subsections with a different number of alternatives, we compute the asymptote parameters per subsection and then compute an overall asymptote using a weighted average in which the weights are proportional to the number of examples in each subsection. and 0 otherwise, except for TruthfulQA, which we empirically compute the first percentile of the scores coming from the full Open LLM Leaderboard and fix the lower asymptote to that value.
@@
Figure 1 gives the results for the first set of experiments. It depicts the average mean absolute error of all methods when predicting LLM benchmark performance, which is measured in percentage points. It shows the competitiveness of Sloth in terms of prediction quality. One important thing to notice is that Sloth errors are similar or lower than the “Size and Tokens” variant, suggesting that the assumed low-dimensional structure between benchmarks results is a good inductive bias. We highlight that the analysis includes recent families like LLaMa 3, Qwen 2, and Yi 1.5. For more details on the tested models and extra related results, including model-specific results, please check Appendix G.1. The extra results are qualitatively similar to the ones in Figure 1, in which Sloth often beats the baselines.
In the second set of experiments, we consider the case in which the two smallest models of the test family are observed at training time. In this way, we can fit (2.2) making both parameters family dependent. For this analysis, we modify all methods that depend on FLOPs, by letting the slope (in addition to the intercept) depend on the family. All variations of our method are kept unchanged. Figure 2 shows the average mean-absolute-error (MAE) across families and benchmarks for different methods. We see that the different variants of our approach are still the most successful ones in performance prediction. We also replicate the observation made by Ruan et al. (2024) that fitting both intercept and slope per family performs poorly. For more detailed results, including family-specific results see Appendix G.2.
@@
4.3 Interpreting the latent skills
In this section, we use the intersection between the two leaderboards, aiming to get more insights from the combined data. Since we have an identifiability result for the “basic” version of Sloth, in which we fix the lower asymptotes $\gamma_{j}$’s and the link function to be sigmoid (see Section 3.2), we opt for interpreting that version of the model. We set $d=3$ as that model version achieved the best prediction results in Figure 11. Figure 4 illustrates the model loadings, $\Lambda$, from which we assign names to the three dimensions based on our subjective interpretation. We include the loadings for $d=2$ and $d=4$ in Appendix H. To complement our exploration, we include Figure 5, which gives us the level curves of different skills (disregarding the family-specific intercept term), and Figure 6 that compare the skills of base and instruction-tuned models; in this figure, we include LLM families with more number of models. In both figures, the numbers are given in terms of standard deviations as the skills are standardized to have zero mean and unitary standard deviation.
Reasoning skill The first dimension, with strong loadings from benchmarks such as GSM8K, MATH, GPQA, MMLU(-PRO), BBH, and MUSR, is labeled “Reasoning.” The benchmarks GSM8k and MATH Lvl 5 consist entirely of mathematical word problems while MMLU/MMLU-PRO and GPQA also contain mathematical and advanced science questions. On the other hand, BBH includes logic deduction and linguistic reasoning puzzles. The strong dependence of BBH on the “Reasoning” skill suggests that in language models, there is an association between logical reasoning, general linguistic ability, and mathematical ability. Finally, MuSR is a benchmark that evaluates “multistep soft reasoning tasks specified in a natural language narrative” (Sprague et al., 2023). Figure 5 shows that Reasoning is primarily a function of model size, with a small dependence on the number of training tokens used. Moreover, the first plot of Figure 6 compares base models versus their instruction-tuned versions in terms of Reasoning and we found that there is no clear rule: instruction tuning can either increase or decrease the ability of an LLM to reason. These findings are robust for different values of $d$ as we can see in the figures of Appendix H.
Knowledge skill The second dimension is positively loaded on ARC, HellaSwag, and Winogrande. These three benchmarks measure the ability of LLMs to remember common sense and basic knowledge; we denominate this skill as “Knowledge”. More specifically, ARC consists of grade school-level science questions, HellaSwag is meant for sentence completion for common scenarios, and Winogrande common sense pronoun resolution problems. Contrasting with Reasoning, Figure 5 shows that Knowledge is highly influenced by both model size and number of training tokens. Moreover, we can see that the range of standard deviations in the middle plot is much greater than in the other two plots, giving us evidence that this skill might be more sensitive to increases in compute resources and less dependent on the LLM families themselves. On the other hand, Figure 6 does not show any strong evidence of the effect of instruction tuning on this skill. These findings are similar to the ones reported in Appendix H for different values of $d$, even though there is no clear one-to-one correspondence of Knowledge in those results.
Instruction following skill: IFEval, which is positively and heavily loaded in this skill, measures how well language models produce answers that follow a verifiable set of instructions; for example, including a keyword $x$ number of times in responses. Therefore, we call it “Instruction Following”. An interesting fact is that instruction tuning has a strong positive effect on this skill for all depicted families we can see in Figure 6. The effect can also be observed in Figure 27 of the appendix. When $d=2$, instruction following gets mixed with other skills and we are not able to see this effect. Regarding Figure 5, we see that Instruction Following depends on both model size and tokens. Unfortunately, this interpretation does not hold when $d=4$ as seen in Appendix H; in that case, instruction following abilities are almost constant with respect to size and number of tokens for models trained on bigger datasets and size is negatively correlated to it when the models are trained on smaller datasets.
@@ 4.4 Predicting LLM performance on downstream tasks
Another useful application of Sloth, which is inspired by Ruan et al. (2024), is to predict the performance in a downstream task for a large model from a relatively small number of prior performance observations from that task. We use Sloth to estimate the latent skills of hypothetical LLMs and then use them to predict the performance of those LLMs in downstream tasks. With this approach, we expand on the experiments of Ruan et al. (2024), which do not consider performance prediction of hypothetical LLMs; as we have seen in Section 4.2 (from the “PCA + FLOPs” baselines in Figures 1 and 2), their method could be adapted to this task but it has, in general, poor predictive performance.
@@
We evaluate this pipeline on two downstream tasks, predicting the performance of meta-llama-3-70B and meta-llama-3-70B-instruct on code completion and meta-llama-3-70B-instruct on emotional intelligence tasks. We fit the same model shown in Section 4.3, but do not include meta-llama-3-70B or meta-llama-3-70B-instruct in the training set (see Figure 30 for the loadings of the latent skills, which is similar to Figure 4). Next, using either HumanEval (Chen et al., 2021) or EQ bench data (Paech, 2024), we fit a regression model with logistic link using latent skills as features and performance on the downstream task as target. Together, this provides us with sufficient information to predict the performance of the held-out models on both tasks with decent accuracy. Figure 7 depicts this logistic curve and the actual values. Moreover, we can see that “Reasoning” is by far the most important skill in predicting coding ability while a mixture of “Reasoning” and “Knowledge” is needed for emotional intelligence (see Figure 30 for a more accurate interpretation of the loadings).
(a) Code Completion
(b) Emotional Intelligence
Figure 7: Predicting model performance in complex downstream tasks like code completion and EQ for LLaMa 3 70B (base/instruct). In the first step, we fit Sloth without including LLaMa 3 70B (base/instruct) in the training set. In the second step, we fit a regression model connecting skills and downstream performance. Finally, we predict LLaMa 3 70B (base/instruct) performance from their predicted Sloth skills.
In Appendix I, a similar test is provided for agentic capability measured by AgentBench (Liu et al., 2023), although to avoid overfitting, in this case, we must fit Sloth with no family-specific intercept.
@@ Appendix A Limitations
Appendix B Identifiability theorem proof
Theorem 3.2.
@@ Appendix B Identifiability theorem proof
Appendix C Connections with factor analysis
@@
Appendix D Motivating the interaction term in Sloth
@@
Appendix E PCA approach formulation
@@
Appendix F Models in our dataset
@@
Appendix G Extra performance prediction results
@@
G.1 Test families have exactly one model in the training set
G.1.1 Average prediction loss across models
@@
G.1.2 Family-specific prediction losses
@@
G.2 Test families have exactly two models in the training set
G.2.1 Average prediction loss across models
@@
Figure 15: The figure shows the average (across LLM families) mean-absolute-error (MAE) (within a family) for different methods. This is a complete version of Figure 2.
@@
G.2.2 Family-specific prediction losses
@@
Appendix H Extra interpretability results
H.1 Results for $d=2$
@@ H.3 Results for $d=4$
Appendix I Extra downstream task plots
(a) Predictions
(b) Parameter Interpretation
@@ Appendix I Extra downstream task plots
Appendix J Insights from the different link functions
In this section, we visually compare Sloth considering trainable and logistic link function $\sigma$, Owen (2024)’s model (“FLOPs (shared intercept)”) and our adaptation of Ruan et al. (2024)’s observational scaling law (“PCA + FLOPs”) described in Appendix E. For this experiment, we study the two Open LLM Leaderboards separately and consider LLaMa-3 and Yi-1.5 families as the test families; we make this choice because both families are popular ones and the training set size is the same for all models in each family, making comparison between models possible (in the x-axis, we use model size). For LLaMA-3, we just include one model from that family in the training set and do not train a family-specific slope for PCA+FLOPs. For Yi-1.5, we include two models in the training set and train a family-specific slope for PCA+FLOPs. In summary, we see that: (i) training the link function can produce a much more flexible scaling law that can better predict performance saturation (e.g., the performance of Yi-1.5 in ARC, HellaSwag etc.), (ii) training no family-specific parameters at all (“FLOPs (shared intercept)”) usually produce poor prediction results, and (iii) PCA+FLOPs often produces flatter curves that underestimate the performance of bigger models, e.g., see Yi-1.5 in TruthfulQA, GSM8k, and MMLU.
@@
Appendix K Comparing against Ruan et al. (2024) in their observational scaling law setting
In this section, we compare Sloth with Ruan et al. (2024)’s observational scaling law; that is, we extract abstract skills using a set of benchmark scores and then use those skills to predict the performance of models of interest in a target downstream task. For this experiment, we use the same data and tasks explored in Section 4.4. For our method, we fit Sloth using benchmark data from all models, including performance data of LLaMa-3-70B models, and extract the skills of each model. For Ruan et al. (2024)’s method, we fit PCA on the benchmark data to extract the skills. For both methods, we set $d=3$ and then fit a regression model with a logistic link to predict downstream performance from skills. Figures 35 and 36 present the prediction results for both methods and Figures 37 and 38 give the loading of both approaches. In both plots, out-sample prediction has a similar prediction error. At the same time, the in-sample fit is better for Sloth in the coding task and for Ruan et al. (2024)’s observational scaling law in the emotional intelligence task. Regarding the loading, it is possible to draw some similarities, e.g., the presence of instruction following skill, but there is no one-to-one correspondence between skills.
(a) Sloth
(b) Obs. Scaling Law (Ruan et al., 2024)
Figure 35: Predicting code completion of LLaMa 3 70B (base/instruct).
(a) Sloth
(b) Obs. Scaling Law (Ruan et al., 2024)
Figure 36: Predicting emotional intelligence of LLaMa 3 70B (base/instruct).

### v4

@@ 4.2 Comparing scaling laws in terms of prediction errors
In the first set of experiments, we consider the case in which only the smallest model of the test family is observed at training time. Because of that reason, we cannot fit the general scaling law in (2.2) in which both the intercept and slope are family dependent. In this scenario, we consider our main baselines to be (i) the model in (2.2) with shared intercept and slope (Owen, 2024) (“FLOPs (shared intercept)”), (ii) the same model but only with shared slope (“FLOPs”), (iii) a version of the PCA idea [^2] We include more details about this approach in Appendix F. proposed by Ruan et al. (2024) in which we predict the principal components using the FLOPs model with shared slope that are then mapped to the benchmark scores (“PCA + FLOPs”), (iv) and our model with trainable activation function but assuming $\Lambda$ is identity (“Size and Tokens”; implies $d=J$). Moreover, we include two versions of Sloth. In the “basic” one, we fix $\sigma$ to be sigmoid, and $\gamma_{j}$’s are given by the $100\%$ over the number of alternatives in the case of multiple-choice benchmarks [^3] When the benchmark has subsections with a different number of alternatives, we compute the asymptote parameters per subsection and then compute an overall asymptote using a weighted average in which the weights are proportional to the number of examples in each subsection. and 0 otherwise, except for TruthfulQA, which we compute the first percentile of the scores coming from the full Open LLM Leaderboard and fix the lower asymptote to that value.
@@
Figure 1 gives the results for the first set of experiments. It depicts the average mean absolute error of all methods when predicting LLM benchmark performance, which is measured in percentage points. It shows the competitiveness of Sloth in terms of prediction quality. One important thing to notice is that Sloth errors are similar or lower than the “Size and Tokens” variant, suggesting that the assumed low-dimensional structure between benchmarks results is a good inductive bias. We highlight that the analysis includes recent families like LLaMa 3, Qwen 2, and Yi 1.5. For more details on the tested models and extra related results, including model-specific results, please check Appendix H.1. The extra results are qualitatively similar to the ones in Figure 1, in which Sloth often beats the baselines. In a second set of experiments, we consider the case in which the two smallest models of the test family are observed at training time. In this way, we can fit (2.2) making both parameters family dependent. The results are qualitatively similar to the one presented on Figure 1 and are presented in Appendix H.2.
@@
4.3 Interpreting the latent skills
In this section, we use the intersection between the two leaderboards, aiming to get more insights from the combined data. Since we have an identifiability result for the “basic” version of Sloth, in which we fix the lower asymptotes $\gamma_{j}$’s and the link function to be sigmoid (see Section B), we opt for interpreting that version of the model. We set $d=3$ as that model version achieved the best prediction results in Figure 11. Figure 2 illustrates the model loadings, $\Lambda$, from which we assign names to the three dimensions based on our subjective interpretation. We include the loadings for $d=2$ and $d=4$ in Appendix I. To complement our exploration, we include Figure 4, which gives us the level curves of different skills (disregarding the family-specific intercept term), and Figure 5 that compare the skills of base and instruction-tuned models; in this figure, we include LLM families with more number of models. In both figures, the numbers are given in terms of standard deviations as the skills are standardized to have zero mean and unitary standard deviation.
Reasoning skill: The first dimension, with strong loadings from benchmarks such as GSM8K, MATH, GPQA, MMLU(-PRO), BBH, and MUSR, is labeled “Reasoning.” The benchmarks GSM8k and MATH Lvl 5 consist entirely of mathematical word problems while MMLU/MMLU-PRO and GPQA also contain mathematical and advanced science questions. On the other hand, BBH includes logic deduction and linguistic reasoning puzzles. The strong dependence of BBH on the “Reasoning” skill suggests that in language models, there is an association between logical reasoning, general linguistic ability, and mathematical ability. Finally, MuSR is a benchmark that evaluates “multistep soft reasoning tasks specified in a natural language narrative” (Sprague et al., 2023). Figure 4 shows that Reasoning is primarily a function of model size, with a small dependence on the number of training tokens used. Moreover, the first plot of Figure 5 compares base models versus their instruction-tuned versions in terms of Reasoning and we found that there is no clear rule: instruction tuning can either increase or decrease the ability of an LLM to reason. These findings are robust for different values of $d$ as we can see in the figures of Appendix I.
Knowledge skill: The second dimension is positively loaded on ARC, HellaSwag, and Winogrande. These three benchmarks measure the ability of LLMs to remember common sense and basic knowledge; we denominate this skill as “Knowledge”. More specifically, ARC consists of grade school-level science questions, HellaSwag is meant for sentence completion for common scenarios, and Winogrande common sense pronoun resolution problems. Contrasting with Reasoning, Figure 4 shows that Knowledge is highly influenced by both model size and number of training tokens. Moreover, we can see that the range of standard deviations in the middle plot is much greater than in the other two plots, giving us evidence that this skill might be more sensitive to increases in compute resources and less dependent on the LLM families themselves. On the other hand, Figure 5 does not show any strong evidence of the effect of instruction tuning on this skill. These findings are similar to the ones reported in Appendix I for different values of $d$.
Instruction following skill: IFEval, which is positively and heavily loaded in this skill, measures how well language models produce answers that follow a verifiable set of instructions; for example, including a keyword $x$ number of times in responses. Therefore, we call it “Instruction Following”. An interesting fact is that instruction tuning has a strong positive effect on this skill for all depicted families we can see in Figure 5. The effect can also be observed in Figure 27 of the appendix. When $d=2$, instruction following gets mixed with other skills and we are not able to see this effect. Regarding Figure 4, we see that Instruction Following depends on both model size and tokens. Unfortunately, this interpretation does not hold when $d=4$ as seen in Appendix I.
Figure 6: Predicting model performance in complex downstream tasks like code completion (“HE”) for LLaMa 3 70B (base/instruct). In the first step, we fit Sloth without including LLaMa 3 70B (base/instruct) in the training set. In the second step, we fit a regression model connecting skills and downstream performance. Finally, we predict LLaMa 3 70B (base/instruct) performance from their predicted Sloth skills.
@@ 4.4 Predicting LLM performance on downstream tasks
Another useful application of Sloth, which is inspired by Ruan et al. (2024), is to predict the performance in a downstream task for a large model from a relatively small number of prior performance observations from that task. We use Sloth to estimate the latent skills of hypothetical LLMs and then use them to predict the performance of those LLMs in downstream tasks. With this approach, we expand on the experiments of Ruan et al. (2024), which do not consider performance prediction of hypothetical LLMs.
@@
We evaluate this pipeline on two downstream tasks, predicting the performance of meta-llama-3-70B and meta-llama-3-70B-instruct on code completion and meta-llama-3-70B-instruct on emotional intelligence tasks. We fit the same model shown in Section 4.3, but do not include meta-llama-3-70B or meta-llama-3-70B-instruct in the training set (see Figure 31 for the loadings of the latent skills, which is similar to Figure 2). Next, using either HumanEval (Chen et al., 2021) or EQ bench data (Paech, 2024), we fit a regression model with logistic link using latent skills as features and performance on the downstream task as target. Together, this provides us with sufficient information to predict the performance of the held-out models on both tasks with decent accuracy. Figure 6 depicts this logistic curve and the actual values for HumanEval; EQ bench results can be found in Appendix J. Moreover, we can see that “Reasoning” is by far the most important skill in predicting coding ability while a mixture of “Reasoning” and “Knowledge” is needed for emotional intelligence (see Figure 31 for a more accurate interpretation of the loadings).
In Appendix J, a similar test is provided for agentic capability measured by AgentBench (Liu et al., 2023), although to avoid overfitting, in this case, we must fit Sloth with no family-specific intercept.
4.5 Predicting performance behavior with scaled inference compute
In this section, we demonstrate how our method, Sloth, combined with concepts like Item Response Theory (IRT) (Reckase, 2009), can predict how the performance of an LLM scales with increased inference through repeated sampling (Brown et al., 2024). For this experiment, we utilize the MATH dataset (Hendrycks et al., 2021), made available by Brown et al. (2024), which evaluates 10 LLMs also included in our dataset. These models are part of the LLaMa 3 Instruct, Gemma, and Pythia families. The process is:
  1. 1.
    Train Sloth on our full dataset of 12 benchmarks, excluding the largest models in the LLaMa 3 Instruct, Gemma, and Pythia families.
  2. 2.
    Fit a logistic regression model for each MATH question in Brown et al. (2024)’s data, using the skills $\theta_{i}(s,t)$ of the 7 training LLMs as covariates to predict the probability of correctly solving those problems.
  3. 3.
    Estimate the skills for the three test models using our scaling law, then predict their probabilities of answering each MATH question correctly via the logistic regressions, and then predict the pass@$k$ metrics for the 3 test models using those probabilities. The predicted pass@$k$ metric for a certain LLM is given by the average of $1-(1-\hat{p}_{j})^{k}$’s (across $j$’s) if $\hat{p}_{j}$ is the predicted probability of the LLM of interest getting question $j$ correct.
Figure 7 illustrates that Sloth can accurately predict test-time scaling behavior for these models. Note that unlike scaling law in Brown et al. (2024), Sloth can predict inference compute gains for hypothetical LLMs before committing resources to training them. This highlights a practical application where practitioners can estimate the potential performance of a hypothetical model at test time given a computational budget.
Figure 7: Sloth can accurately predict test-time scaling behavior for LLaMa 3 Instruct, Gemma, and Pythia models when coupled with Item Response Theory (IRT).
@@ Appendix A Limitations
Appendix B Identifiability of model parameters
Assumption B.1 (Identifiability constraints).
Assume that
$\frac{1}{n}\sum_{i=1}^{n}\theta^{(i)}=0,\quad\frac{1}{n}\sum_{i=1}^{n}{\theta^{(i)}}^{\top}\theta^{(i)}=\Psi,$
and that $\text{rank}(\Lambda)=d$ and $\text{rank}(X)=p$, where $\theta^{(i)}$ denotes the $i$-th row of $XB$ and $\Psi$ is a positive definite matrix.
One possible choice for the covariance matrix is $\Psi=I_{d}$ (Chen et al., 2019), which assumes uncorrelated skills. One implicit implication of Assumption B.1 is that $n\geq p\geq d$ must be satisfied, otherwise the covariance matrix cannot be full rank. This condition is satisfied in our experiments. Under Assumption B.1, we show the identifiability of the model parameters up to a transformation of $\Lambda$ tied to a transformation of $B$, which leaves the outputs of the model unchanged. This means that we can potentially approximate the true values for $\Lambda$ and $B$ up to a transformation, which is usually the norm within the class of exploratory factor analysis models.
Theorem B.2.
then, under the Assumption B.1, we have $\tilde{b}=b$, $\tilde{\Lambda}=\Lambda M$, and $\tilde{B}=B(M^{\top})^{-1}$ for an invertible matrix $M\in{\mathbf{R}}^{d\times d}$. In particular, $M$ is orthogonal if $\Psi=I_{d}$, i.e., $M^{\top}M=MM^{\top}=I_{d}$.
Proof.
@@ Appendix B Identifiability of model parameters
From our proof, we can see that the matrix $M$ is dependent on the specification of $\Psi$.
Appendix C Model fitting
[…]
Appendix D Connections with factor analysis
@@
Appendix E Motivating the interaction term in Sloth
@@
Appendix F PCA approach formulation
@@
Appendix G Models in our dataset
@@
Appendix H Extra performance prediction results
@@
H.1 Test families have exactly one model in the training set
H.1.1 Average prediction loss across models
@@
H.1.2 Family-specific prediction losses
@@
H.2 Test families have exactly two models in the training set
H.2.1 Average prediction loss across models
@@
Figure 15: The figure shows the average (across LLM families) mean-absolute-error (MAE) (within a family) for different methods.
@@
H.2.2 Family-specific prediction losses
@@
Appendix I Extra interpretability results
I.1 Results for $d=2$
@@ I.3 Results for $d=4$
Appendix J Extra downstream task plots
J.1 Emotional Intelligence
Figure 29: Predicting model performance in complex downstream tasks like emotional intelligence (“EQ”) for LLaMa 3 70B (base/instruct). In the first step, we fit Sloth without including LLaMa 3 70B (base/instruct) in the training set. In the second step, we fit a regression model connecting skills and downstream performance. Finally, we predict LLaMa 3 70B (instruct) performance from their predicted Sloth skills.
J.2 Agentic Capabilities
Refer to caption Refer to caption
@@
Appendix K Insights from the different link functions
In this section, we visually compare Sloth considering trainable and logistic link function $\sigma$, (Owen, 2024)’s model (“FLOPs (shared intercept)”) and our adaptation of Ruan et al. (2024)’s observational scaling law (“PCA + FLOPs”) described in Appendix F. For this experiment, we study the two Open LLM Leaderboards separately and consider LLaMa-3 and Yi-1.5 families as the test families; we make this choice because both families are popular ones and the training set size is the same for all models in each family, making comparison between models possible (in the x-axis, we use model size). For LLaMA-3, we just include one model from that family in the training set and do not train a family-specific slope for PCA+FLOPs. For Yi-1.5, we include two models in the training set and train a family-specific slope for PCA+FLOPs. In summary, we see that: (i) training the link function can produce a much more flexible scaling law that can better predict performance saturation (e.g., the performance of Yi-1.5 in ARC, HellaSwag etc.), (ii) training no family-specific parameters at all (“FLOPs (shared intercept)”) usually produce poor prediction results, and (iii) PCA+FLOPs often produces flatter curves that underestimate the performance of bigger models, e.g., see Yi-1.5 in TruthfulQA, GSM8k, and MMLU.
@@
Appendix L Comparing against Ruan et al. (2024) in their observational scaling law setting
In this section, we compare Sloth with Ruan et al. (2024)’s observational scaling law; that is, we extract abstract skills using a set of benchmark scores and then use those skills to predict the performance of models of interest in a target downstream task. For this experiment, we use the same data and tasks explored in Section 4.4. For our method, we fit Sloth using benchmark data from all models, including performance data of LLaMa-3-70B models, and extract the skills of each model. For Ruan et al. (2024)’s method, we fit PCA on the benchmark data to extract the skills. For both methods, we set $d=3$ and then fit a regression model with a logistic link to predict downstream performance from skills. Figures 36 and 37 present the prediction results for both methods and Figures 38 and 39 give the loading of both approaches. In both plots, out-sample prediction has a similar prediction error. At the same time, the in-sample fit is better for Sloth in the coding task and for Ruan et al. (2024)’s observational scaling law in the emotional intelligence task. Regarding the loading, it is possible to draw some similarities, e.g., the presence of instruction following skill, but there is no one-to-one correspondence between skills.
Refer to caption Refer to caption
Figure 36: Predicting code completion of LLaMa 3 70B (base/instruct) with Sloth vs Obs. Scaling Law (Ruan et al., 2024).
Refer to caption Refer to caption
Figure 37: Predicting emotional intelligence of LLaMa 3 70B (base/instruct) with Sloth vs Obs. Scaling Law (Ruan et al., 2024).